### v1

@@ A.3.1 First order terms
A.3.2 Second order terms
@@ Appendix C Additional results
Table C.2: Estimated correlation matrices of the items in the neuroticism scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Table C.3: Estimated correlation matrices of the items in the extroversion scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Table C.4: Estimated correlation matrices of the items in the conscientiousness scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=1.6$ (bottom panel). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
1 9,814,457,557.73 16,011.33 11.82 1.14 0.65
2 2,424.07 10.07 0.80 0.90 1.42
3 5.48 0.65 0.99 0.80 77.14
4 0.88 0.92 0.61 12.66 222,528.08
5 0.89 0.88 36.01 55,420.33 995,017,243,197.60
Table C.5: Decentralized Pearson residual, $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})$, of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of a neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=1.6$.
Figure C.5: Dot plot of cells for the neuroticism item adjective pair “envious” and “not envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s DPR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})$, at robust parameter estimate with tuning constant $c=1.6$. Note that some cells have an extremely large DPR (substantially larger than ideal value 1; see also Table C.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with DPR $\geq 4$, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.

### v2

@@ A.2.1 First order terms
A.2.2 Second order terms
@@ Appendix C Additional results
Table C.2: Estimated correlation matrices of the items in the neuroticism scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “calm” (N1_P), “angry” (N1_N), “relaxed” (N2_P), “tense” (N2_N), “at ease” (N3_P), “nervous” (N3_N), “not envious” (N4_P), “envious” (N4_N), “stable” (N5_P), “unstable” (N5_N), “contented” (N6_P), and “discontented” (N6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Table C.3: Estimated correlation matrices of the items in the extroversion scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “extraverted” (E1_P), “introverted” (E1_N), “energetic” (E2_P), “unenergetic” (E2_N), “talkative” (E3_P), “silent” (E3_N), “bold” (E4_P), “timid” (E4_N), “assertive” (E5_P), “unassertive” (E5_N), “adventurous” (E6_P), and “unadventurous” (E6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
Table C.4: Estimated correlation matrices of the items in the conscientiousness scale from the data in Arias et al., (2020, Sample 1; $N=725$) using MLE (top panel) and our robust estimator with tuning constant $c=0.6$ (bottom panel). The items are “calm” (C1_P), “angry” (C1_N), “relaxed” (C2_P), “tense” (C2_N), “at ease” (C3_P), “nervous” (C3_N), “not envious” (C4_P), “envious” (C4_N), “stable” (C5_P), “unstable” (C5_N), “contented” (C6_P), and “discontented” (C6_N). For the item naming given in parentheses, items with identical identifier (the integer after the first “N”) are polar opposites, where a last character “P” refers to the positive opposite and “N” to the negative opposite.
@@
1 9,814,457,556.73 16,010.33 10.82 0.14 $-0.35$
2 2,423.07 9.07 $-0.20$ $-0.10$ 0.42
3 4.48 $-0.35$ $-0.01$ $-0.20$ 76.14
4 $-0.12$ $-0.08$ $-0.39$ 10.66 222,527.08
5 $-0.11$ $-0.12$ 35.01 55,419.33 995,017,243,196.60
Table C.5: Pearson residual, $\widehat{f}_{N}(x,y)\big{/}p_{xy}(\widehat{\bm{\theta}}_{N})-1$, of each response $(x,y)$ for the “not envious” ($X$)–“envious” ($Y$) item pair in the measurements of Arias et al., (2020) of a neuroticism scale. Estimate $\widehat{\bm{\theta}}_{N}$ was computed with tuning constant $c=0.6$.
Figure C.5: Dot plot of cells for the neuroticism item adjective pair “envious” and “not envious” in the data of Arias et al., (2020), where each item has five Likert-type response options, anchored by “very inaccurate” (= 1) and “very accurate” (= 5). Each dot’s size is proportional to the relative empirical frequency of its associated contingency table cell, $\widehat{f}_{N}(x,y)$, whereas its color varies by the value of the cell’s PR, $\widehat{f}_{N}(x,y)/p_{xy}(\widehat{\bm{\theta}}_{N})-1$, at robust parameter estimate with tuning constant $c=0.6$. Note that some cells have an extremely large PR (substantially larger than ideal value 0; see also Table C.5), indicating a poor model fit for those cells. We therefore censored the color scale so that a fully saturated red is assigned to all cells with PR $\geq 3$, i.e., the empirical frequency $\widehat{f}_{N}(x,y)$ being at least four times as large as the model frequency $p_{xy}(\widehat{\bm{\theta}}_{N})$.